Circular No. 7697

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2001 Q5 (LINEAR-NEAT) K. J. Lawrence, E. F. Helin, and S. H. Pravdo, Jet Propulsion Labo-

ratory, reported the discovery by the NEAT program of a comet, having a nuclear condensation of diameter ∼ 6′′ and a 100′′ tail toward the southwest, on CCD images taken with the 1.2-m Schmidt telescope at Palomar on Aug.

28. Following posting on the NEO Confirmation Page, T. B. Spahr (Minor Planet Center) identified this object with an object reported as asteroidal by the LINEAR program (first detected on Aug. 17 at m_{2} = 18.6–19.4). Other reported physical descriptions from CCD images include: Aug. 29.04 UT, m_{1} = 16.0, tail > 50′′ long in p.a. 237^{o} (M. Tichý, Kleť, 0.57-m reflector); 29.05, small coma, tail at least 4′ long (L. Šarounová, Ondřejov, 0.65-m reflector); 29.08, diffuse (A. Galád and D. Kalmančok, Modra, 0.6m reflector); 29.30, 40′′ tail in p.a. 243^{o} (K. Smalley, Louisburg, KS, 0.75-m reflector); 29.38, well-condensed coma, broad tail 2′ long in p.a. 246^{o} (D. Balam, Dominion Astrophysical Observatory, 1.82-m Plaskett telescope); 29.43, m_{1} = 16.9, tail ∼ 70′′ long in p.a. ∼ 240^{o} (P. J. Shelus, McDonald Observatory, 0.76-m reflector).

2001 UT α_{2000} δ_{2000} m_{1} Observer
Aug. 17.40174 2^{h}02^{m}03.^{s}23 +20^{o}55′57′′.0  LINEAR
28.35137 2 07 35.68 +22 58 30.2 17.3 NEAT

Additional astrometry and the following orbital elements by B. G. Marsden (from 35 observations, Aug. 17–29) appear on MPEC 2001-Q69:

T = 2001 June 11.7358 TT  ω = 6.^{o}4144
e = 0.419096  Ω = 336.3234 }2000.0
q = 2.044589 AU  i = 10.9591
a = 3.519665 AU n^{o} = 0.1492629 P = 6.603 years
2001 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Aug. 29 2^{h}07.^{m}81 +23^{o}05.^{′}3 1.467 2.132 117.^{o}6 24.^{o}8 16.1
Sept. 8 2 09.42 +24 42.8 1.402 2.155 126.0 22.2 16.1
18 2 07.70 +26 03.5 1.349 2.180 135.1 19.0 16.0
28 2 02.74 +27 02.7 1.312 2.206 144.5 15.3 16.0
Oct. 8 1 55.18 +27 36.6 1.295 2.235 153.9 11.4 16.1
18 1 46.08 +27 43.8 1.299 2.266 161.5  8.0 16.1
28 1 36.88 +27 26.5 1.327 2.298 163.7  7.0 16.2
2001 August 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green